# Concepts in Motion: Temporal Bottlenecks for Interpretable Video Classification

### v1

@@ 5 Discussion
MoTIF demonstrates that the CBM principle can be extended effectively to video data. Our framework outperforms zero-shot classification and a linear probe that averages over windows, proving that the architecture yields not only interpretability but also improved predictive performance. MoTIF reaches competitive accuracy levels despite its explicit interpretability constraints. Notably, MoTIF outperforms non-interpretable baselines (works that report results) on the Breakfast dataset, which contains the longest video sequences, and provides fine-grained insights by identifying the ordered steps in tasks such as meal preparation.
MoTIF inherits the flexibility of CBMs. In this work, we focus on CLIP-based variants and demonstrate that CLIP backbones trained on images remain effective when applied to video classification. We further show that adapted CLIP backbones, such as SigLIP or PE, achieve superior performance compared to their counterparts while maintaining the same parameter complexity. Especially PE that has been adapted to embed videos. Future studies could explore backbones tailored for video embeddings (Wang et al., 2024b), which must not necessarily be text-image aligned, e.g., follow the approach of (Prasse et al., 2025) or (Lee et al., 2025). Moreover, the per-channel temporal self-attention block opens the possibility of applying MoTIF beyond video to other modalities with temporal structure, such as time series or text, or even other applications where channel preserving is important.
Interpretability remains central. MoTIF’s three-level explanation property preserves concept representations without the entanglement issues that arise in standard attention mechanisms. Therefore, MoTIF gives valuable insights into the compositional structure of concepts the model has learned for its classification. However, the multiplication of concept activations with attention weights can sometimes yield unexpected outcomes, which complicates a direct comparison with standard action recognition benchmarks. Future adaptations that disentangle this effect may improve comparability.
Several open challenges remain. Selecting window size is non-trivial since actions span variable durations and can be missed if windows are too short, too long, or misaligned with action speed. Similarly, extracting richer and more action-relevant concepts directly from videos could improve performance, particularly for abstract datasets such as SSv2. In MoTIF, however, our objective was not to optimize for the most suitable concept set, but rather to demonstrate the functionality of the proposed architecture. One could also investigate how to model dependencies between concepts while maintaining interpretability. Capturing interactions is key for complex temporal reasoning, but care must be taken not to collapse into uninterpretable feature mixtures.
@@ 6 Conclusion
MoTIF, to the best of our knowledge, is the first concept bottleneck framework tailored to video data that identifies when and which concept is activated with diagonal attention. Through extensive evaluation across multiple embedding backbones and diverse datasets, we demonstrated its effectiveness and generality. Beyond strong empirical performance, MoTIF opens a new research direction by enabling fine-grained explanations of which visual motifs drive predictions in dynamic video settings. We believe this work can stimulate further exploration of concept-based models for video understanding, bridging the gap between interpretability and high-capacity temporal architectures.
Acknowledgments
This research was supported in part by the German Federal Ministry for Economic Affairs and Climate Action of Germany (BMWK), and in part by the German Federal Ministry for Research, Technology, and Space (BMFTR).
@@ A.1 Hyperparameter Settings
  • Architecture. One Transformer layer with per-channel (diagonal) attention; classifier weights constrained to be nonnegative. As stated in Section 2, the diagonal attention faces the trade-off: interpretability through concept isolation versus representational power. Thus, depth (stacking of more than one per-channel temporal block) mainly repeats the same constrained operation, yielding limited or no performance gains.
@@
Dataset-specific deviations: for SSv2 and UCF101 we use learning rate $0^{-4}, 1\ell_{1}$ penalty $10^{-4}$, and sparsity penalty $10^{-4}$; for SSv2, the non-negativity constraint in AdamW is disabled. HMDB51 uses a shorter window size of 8. For Breakfast, the window size is increased to 32 to account for longer videos.
@@ A.2 Computation Time and Complexity
For short videos (small $T$) epoch times remain nearly unchanged while GPU memory differs substantially. For longer sequences, as in Breakfast, disabling diagonal attention reduces memory use considerably; the runtime and memory gap grows with sequence length since complexity increases with $T$ (see Sec. 2) if $H<<C$. The runtime across backbones is similar in magnitude; for example the best-performing perception encoder in MoTIF required 4.55 s. Video embedding time is excluded because embeddings are reusable across models and vary with the embedding backbone.
@@ B.1 Where motifs help to understand Misclassifications
Although MoTIF does not always predict the correct class, its structure makes the reasoning process transparent by decomposing decisions into concepts and their temporal interactions. In this work, our emphasis was on the architecture rather than on designing the most suitable concept sets. We expect that future advances in concept extraction for video data will further improve performance, complementing MoTIF’s interpretability with stronger predictive accuracy.
@@ Appendix C Additional Ablations
Random seed.
Since most experiments were run with a fixed seed of 42 for reproducibility, we additionally ablate the effect of varying the random seed (39,40,41,42,43) for both MoTIF and the linear probe. As Figure 12 illustrates, the influence of the seed depends on the dataset. For HMDB51 the effect is negligible, whereas for Breakfast — particularly with the RN/50 backbone — we observe noticeably higher variance. Nevertheless, MoTIF consistently outperforms the linear probe. Moreover, the figure indicates that the reported numbers in Table 1 are conservative: even higher scores are achievable, but we deliberately refrain from reporting maxima and instead provide a representative overview.
@@
Additionally, since embeddings depend on the random choice of a representative image per window (previously fixed at seed 42), we re-embedded each dataset with five seeds. For each embedding variant we trained and evaluated MoTIF, the linear probe, and the zero-shot runs using a fixed training seed of 42 to isolate the effect of image-selection randomness. Table 4 reports the mean and standard deviation across embeddings (seeds 39–43). Std values are small overall; the largest observed variability is for Breakfast with RN/50 (std = 1.1).
@@
Appendix D Additional Performance Comparisons
D.1 Comparison to other models
Table 5 provides an extended version of the main paper’s results, including accuracies from non-interpretable baselines. In particular, we compare against TSM (Lin et al., 2019), the No Frame Left Behind model by Liu et al. (2021), and VideoMAE V2 (Wang et al., 2023), all of which report results on our selected datasets. As expected, these models generally outperform our interpretable variant. However, on Breakfast, we exceeded the performance of two of the baselines that report scores. Importantly, our objective is not to surpass state-of-the-art benchmarks, but to demonstrate a novel MoTIF framework for video data that provides unique interpretability insights.
Table 5: Performance comparison (% Top-1 accuracy). MoTIF on Breakfast Actions, HMDB51, UCF101, and SSv2 compared to other video classification models.
@@ D.1 Comparison to other models
Baselines (ours)
MoTIF (RN/50) 52.8 $\pm$ 6.9 62.8 $\pm$ 1.1 82.8 $\pm$ 0.6 16.0
MoTIF (ViT-B/32) 53.4 $\pm$ 6.9 65.3 $\pm$ 1.8 85.6 $\pm$ 1.2 17.5
MoTIF (ViT-L/14) 69.3 $\pm$ 6.2 73.3 $\pm$ 1.0 93.2 $\pm$ 0.7 20.4
MoTIF (SigLIP-L/14) 73.5 $\pm$ 8.6 73.2 $\pm$ 2.4 94.0 $\pm$ 0.8 22.4
MoTIF (PE-L/14) 83.6 $\pm$ 6.5 79.6 $\pm$ 0.3 95.4 $\pm$ 0.7 30.0
Existing video models
TSM (Lin et al., 2019) 59.1111from (Liu et al., 2021) with eight clusters and all frames (cumulative) 73.5 95.9 61.7
No frame left behind (Liu et al., 2021) 62.0111from (Liu et al., 2021) with eight clusters and all frames (cumulative) 73.4111from (Liu et al., 2021) with eight clusters and all frames (cumulative) 96.4111from (Liu et al., 2021) with eight clusters and all frames (cumulative) 62.7111from (Liu et al., 2021) with eight clusters and all frames (cumulative)
@@
D.2 Accuracies Testsplits
Tables 68 report Top-1 accuracies (%) on the test splits. Mean and standard deviation are computed across the shown splits (Breakfast: 4 splits; HMDB51 and UCF: 3 splits). For each backbone we list the MoTIF result (bold), the corresponding linear probe, and the zero-shot baseline. The reported standard deviation quantifies variability between test splits.
Model s1 s2 s3 s4 Mean Std
CLIP RN/50 (MoTIF) 55.6 47.4 47.0 61.2 52.8 6.9
Linear probe (RN/50) 43.3 25.7 32.5 44.4 36.5 9.0
Zero-shot (RN/50) 17.3 19.2 16.0 21.9 18.6 2.6
CLIP B/32 (MoTIF) 56.3 44.9 51.3 61.0 53.4 6.9
Linear probe (B/32) 42.6 27.5 31.6 47.0 37.2 9.1
Zero-shot (B/32) 24.6 19.8 22.0 26.4 23.2 2.9
CLIP L/14 (MoTIF) 71.5 62.3 66.7 76.8 69.3 6.2
Linear probe (L/14) 61.3 44.7 48.7 66.4 55.3 10.2
Zero-shot (L/14) 32.4 28.3 26.5 37.0 31.1 4.7
SigLIP L/14 (MoTIF) 76.1 62.1 73.1 82.7 73.5 8.6
Linear probe (SigLIP L/14) 59.9 44.7 53.3 70.6 57.1 10.9
Zero-shot (SigLIP L/14) 28.9 18.0 20.9 26.6 23.6 5.0
PE L/14 (MoTIF) 87.3 74.7 83.1 89.4 83.6 6.5
Linear probe (PE L/14) 81.0 58.7 72.0 79.9 72.9 10.3
Zero-shot (PE L/14) 40.5 36.6 36.8 51.6 41.4 7.0
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 64.1 62.3 62.1 62.8 1.1
Linear probe (RN/50) 58.6 60.1 59.2 59.3 0.8
Zero-shot (RN/50) 29.3 30.1 30.1 29.8 0.5
CLIP B/32 (MoTIF) 65.9 66.8 63.3 65.3 1.8
Linear probe (B/32) 62.0 63.0 59.8 61.6 1.6
Zero-shot (B/32) 38.4 37.9 38.0 38.1 0.3
CLIP L/14 (MoTIF) 73.8 73.9 72.2 73.3 1.0
Linear probe (L/14) 68.5 68.8 67.9 68.4 0.5
Zero-shot (L/14) 45.8 45.6 45.6 45.7 0.1
SigLIP L/14 (MoTIF) 74.8 74.4 70.4 73.2 2.4
Linear probe (SigLIP L/14) 66.3 66.0 62.6 65.0 2.1
Zero-shot (SigLIP L/14) 48.4 50.0 49.5 49.3 0.8
PE L/14 (MoTIF) 79.9 79.3 79.6 79.6 0.3
Linear probe (PE L/14) 74.0 75.0 74.1 74.4 0.6
Zero-shot (PE L/14) 56.7 57.3 56.2 56.7 0.6
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 82.4 82.5 83.4 82.8 0.6
Linear probe (RN/50) 80.7 80.0 79.3 80.0 0.7
Zero-shot (RN/50) 56.5 56.9 58.3 57.2 0.9
CLIP B/32 (MoTIF) 84.4 85.7 86.7 85.6 1.2
Linear probe (B/32) 82.2 82.5 83.6 82.8 0.7
Zero-shot (B/32) 59.4 60.2 60.1 59.9 0.4
CLIP L/14 (MoTIF) 92.4 93.7 93.6 93.2 0.7
Linear probe (L/14) 88.8 91.0 90.3 90.0 1.1
Zero-shot (L/14) 71.1 70.6 70.1 70.6 0.5
SigLIP L/14 (MoTIF) 93.3 93.8 94.9 94.0 0.8
Linear probe (SigLIP L/14) 90.0 91.0 90.5 90.5 0.5
Zero-shot (SigLIP L/14) 80.0 81.9 79.2 80.4 1.4
PE L/14 (MoTIF) 94.6 95.7 95.8 95.4 0.7
Linear probe (PE L/14) 94.6 93.9 95.0 94.5 0.6
Zero-shot (PE L/14) 73.8 75.6 74.4 74.6 0.9
Appendix E Algorithms

### v2

@@ 5 Discussion
MoTIF introduces a transformer-based concept bottleneck architecture for temporal data. The core architectural contribution of MoTIF is a temporal concept reasoning module that operates on sequences of interpretable concept activations. By combining per-concept temporal self-attention with transformer-style processing, MoTIF models when individual concepts recur and how their temporal patterns contribute to video-level predictions, while preserving explicit concept attribution. Across datasets, this design consistently outperforms zero-shot classification, global CBMs that average over windows, and prior video CBMs, demonstrating that transformer-based temporal reasoning can be integrated into concept bottleneck models without sacrificing interpretability.
Temporal concept quality is decisive for performance. While the MoTIF architecture provides the capacity for temporal reasoning, its effectiveness depends critically on the availability of concepts that reflect actions and interactions over time. Agentic concept discovery substantially improves performance—most notably on SSv2—by supplying object- and action-centric concepts that are temporally expressive. Together, these findings indicate that concept discovery and architectural advances are complementary: MoTIF provides the temporal reasoning backbone, while agent-generated concepts supply the temporally expressive evidence needed to handle challenging datasets.
MoTIF is modular and compatible with modern video backbones. We demonstrate that MoTIF can be paired with a range of embedding backbones, including CLIP and video-adapted Perception Encoders, with consistent gains across architectures. MoTIF-ST extends the framework to space–time transformers, showing that the proposed concept reasoning module is not restricted to a single transformer design. This modularity enables future integration with stronger video representations (Wang et al., 2024b) or alternative concept sources (Prasse et al., 2025; Lee et al., 2025b), while retaining interpretable temporal reasoning.
Interpretability is central to both explanation and model analysis. MoTIF’s three-level explanation interface—global, local, and temporal—exposes how individual concepts contribute across time without the entanglement introduced by standard attention mechanisms. This enables errors to be traced back to missing or ambiguous temporal concepts rather than opaque feature interactions, guiding refinement of the concept set and model behavior while maintaining competitive performance.
Challenges and future work. Selecting an appropriate temporal window size remains an important design consideration, as actions span variable durations and temporal granularity. A key next step for the community is to evaluate adaptive and multi-scale windowing strategies to improve robustness without sacrificing interpretability. Improving concept discovery—particularly for abstract interactions—is a high-impact next step: the substantial gains from agentic concepts on SSv2 indicate that richer concept sets yield larger benefits than architectural scaling alone. Finally, allowing cross-concept interactions increases expressivity but breaks faithful attribution; MoTIF makes this trade-off explicit rather than implicit by enforcing concept-level temporal isolation, while future work may explore controlled relaxations of this constraint.
@@ 6 Conclusion
We introduced MoTIF, a transformer-based concept bottleneck framework for video that combines temporal concept reasoning with agentic concept discovery. MoTIF employs diagonal temporal attention to track individual concepts over time, enabling faithful attribution of when and which concepts drive predictions while exposing a controllable accuracy–interpretability trade-off, where cross-concept attention can recover performance on demanding datasets at the cost of transparency. Our agentic discovery module enables the unsupervised discovery of temporally expressive object- and action-centric concepts directly from video, substantially narrowing the performance gap between interpretable models and black-box video transformers on challenging benchmarks. Overall, MoTIF provides a principled and scalable framework for learning, analyzing, and explaining temporal concept representations, enabling interpretable video understanding with competitive performance.
Acknowledgements
This research was supported in part by the German Federal Ministry for Economic Affairs and Climate Action of Germany (BMWK) and in part by the German Federal Ministry for Research, Technology, and Space (BMFTR), and was partially conducted during an internship at Ramblr.ai.
@@ A.1 Hyperparameter Settings
  • Architecture. One Transformer layer with per-channel (diagonal) attention; classifier weights constrained to be nonnegative. As stated in Section 2, the diagonal attention faces the trade-off: interpretability through concept isolation versus representational power. Thus, additional depth yields diminishing returns under strict concept isolation, motivating systematic study of selectively enabling cross-concept interactions.
@@
Dataset-specific deviations: for SSv2 and UCF101 we use learning rate $0^{-4}, 1\ell_{1}$ penalty $10^{-4}$, and sparsity penalty $10^{-4}$; for SSv2, the non-negativity constraint in AdamW is disabled and batch size 128. HMDB51 uses a shorter window size of 8. For Breakfast, the window size is increased to 32 to account for longer videos.
@@ A.2 Computation Time and Complexity
For short videos (small $T$) epoch times remain nearly unchanged while GPU memory differs substantially. For longer sequences, as in Breakfast, disabling diagonal attention reduces memory use considerably; the runtime and memory gap grows with sequence length since complexity increases with $T$ (see Sec. 2) if $H\ll C$. The runtime across backbones is similar in magnitude; for example the best-performing perception encoder in MoTIF required 4.55 s. Video embedding time is excluded because embeddings are reusable across models and vary with the embedding backbone.
@@ A.3 Backbone Integration
A.4 Space-Time Attention Extension
While the standard MoTIF architecture applies temporal attention independently for each concept, we extend it with a factorized space-time attention mechanism (Bertasius et al., 2021) that enables concept interactions while preserving interpretability.
The CBMTransformerST architecture factorizes attention into two sequential components: spatial attention across concepts at each time step, followed by the original per-channel temporal attention.
PerTimeSpatialBlock. This module computes attention across concepts at each time step independently. Given input features $X\in\mathbb{R}^{B\times T\times C}$, spatial attention produces attention scores $W_{s}\in\mathbb{R}^{B\times T\times C\times C}$ where $W_{s}[b,t,i,j]$ represents the attention weight from concept $i$ to concept $j$ at time step $t$.
Unlike temporal attention, which maintains concept isolation, spatial attention allows concepts to interact within each temporal frame. To preserve concept interpretability while enabling controlled spatial interaction, we employ three identity-preserving mechanisms:
  • Identity bias: We add a bias term $\alpha_{I}=1.0$ to the diagonal of the attention score matrix, encouraging self-attention and reducing cross-concept mixing.
  • Spatial gating: The spatial attention output is scaled by a gating factor $\beta_{s}\in[0,1]$ before being added to the residual connection: $X^{\prime}=X+\beta_{s}\cdot\text{SpatialAttn}(X)$. With $\beta_{s}=0.1$ (default), the spatial mixing contributes 10% to the output, preserving 90% of the original concept identity.
  • Per-channel FFN: The feed-forward network in the spatial block uses per-channel convolutions (grouped by $C$), ensuring no cross-concept mixing occurs in the FFN, matching the design of the temporal block.
These mechanisms ensure that concepts remain separable and interpretable: the spatial mixing is controlled and identity-preserving, the subsequent per-channel temporal attention maintains diagonal structure, and the final concept activations $Z_{t,c}$ and spatial attention maps $W_{s}$ enable attribution to individual concepts while revealing their interactions.
SpaceTimeBlock. The block applies spatial attention first to enable concept interactions at each time step, followed by per-channel temporal attention that maintains diagonal structure.
This factorization reduces computational complexity from $\mathcal{O}(T^{2}C^{2})$ for full space-time attention to $\mathcal{O}(TC^{2}+CT^{2})$, while producing separate interpretable attention maps for spatial ($W_{s}\in\mathbb{R}^{B\times T\times C\times C}$) and temporal ($W_{t}\in\mathbb{R}^{B\times C\times T\times T}$) components.
Exemplary Explanations. MoTIF with CBMTransformerST (MoTIF-ST) applies two sequential attention mechanisms—spatial followed by temporal—yielding two complementary attention matrices for interpretation.
  • The spatial attention block captures concept–concept interactions within each frame. It indicates which concepts influence a given concept and highlights additional concepts that are relevant at that moment. This provides a structured view of how concepts relate to each other spatially.
  • The temporal attention block captures when a concept becomes important. It identifies the time steps at which a specific concept contributes most to the final prediction, thereby revealing the temporal structure of the activity.
Figure 6 shows two examples from Breakfast and UCF101, including their attention matrices and the corresponding explanations.
Figure 6: MoTIF-ST explanations. Correctly classified examples from Breakfast and UCF101, illustrating MoTIF with a space–time transformer (ViT-L/14) and the corresponding temporal and spatial attention matrices.
@@ B.1 Where motifs help to understand Misclassifications
Although MoTIF does not always predict the correct class, its structure makes the reasoning process transparent by decomposing decisions into concepts and their temporal interactions. In this work, our emphasis was on the architecture rather than on designing the most suitable concept sets. Follow-up experiments should evaluate stronger video-specific concept extraction pipelines to improve performance while retaining MoTIF’s interpretability.
@@ B.3 Interpreting Temporal Dependencies
B.4 Synthetic Temporal Data
The synthetic experiment is designed to isolate temporal reasoning by removing all appearance cues. Each “video” is a sequence of length-N (matching the Breakfast distribution), where each frame is a 1024-dimensional vector. We define five temporal classes: (1) ascending, (2) descending, (3) U-shape, (4) inverted-U, and (5) periodic oscillation. Class identity is determined only by the temporal trajectory of 30% of the dimensions; the remaining dimensions are filled with noise. Thus, the task can only be solved by learning the temporal order—static appearance information is uninformative by design.
To examine whether the model can learn interpretable temporal structure, we provide 37 artificial temporal pattern concepts. Each concept represents a prototypical analytic temporal signal (ascending, descending, S-curve, exponential growth/decay, chirp, sawtooth, step-up/down, periodic-fast, oscillating-decay, double-peak, plateau, etc.). All embeddings encode temporal shape only, without semantics.
The training setup matches our real pipelines, allowing a direct test of MoTIF’s temporal inductive biases. MoTIF achieves 86.97% accuracy on the original (ordered) sequences but only 21.06% when frames are randomly permuted (chance 20%). This 4.1× drop shows that MoTIF relies on temporal ordering rather than static frame statistics. An image-level CBM baseline (“Global CBM”), which collapses all frames into a single vector, reaches only 35.5%, confirming that methods without temporal modeling cannot solve this task.
@@ Appendix C Additional Ablations
Random seed. Since most experiments were run with a fixed seed of 42 for reproducibility, we additionally ablate the effect of varying the random seed (39,40,41,42,43) for both MoTIF and the global CBM. As Figure 12 illustrates, the influence of the seed depends on the dataset. For HMDB51 the effect is negligible, whereas for Breakfast — particularly with the RN/50 backbone — we observe noticeably higher variance. Nevertheless, MoTIF consistently outperforms the global CBM. Moreover, the figure indicates that the reported numbers in Table 1 are conservative: even higher scores are achievable, but we deliberately refrain from reporting maxima and instead provide a representative overview.
Additionally, since embeddings depend on the random choice of a representative image per window (previously fixed at seed 42), we re-embedded each dataset with five seeds. For each embedding variant we trained and evaluated MoTIF, the global CBM , and the zero-shot runs using a fixed training seed of 42 to isolate the effect of image-selection randomness. Table 5 reports the mean and standard deviation across embeddings (seeds 39–43). Std values are small overall; the largest observed variability is for Breakfast with RN/50 (std = 1.1).
@@
C.1 Accuracies Testsplits
Tables 68 report Top-1 accuracies (%) on the test splits. Mean and standard deviation are computed across the shown splits (Breakfast: 4 splits; HMDB51 and UCF: 3 splits). For each backbone we list the MoTIF result (bold), the corresponding global CBM, and the zero-shot baseline. The reported standard deviation quantifies variability between test splits.
Model s1 s2 s3 s4 Mean Std
CLIP RN/50 (MoTIF) 55.6 47.4 47.0 61.2 52.8 6.9
Global CBM (RN/50) 43.3 25.7 32.5 44.4 36.5 9.0
Zero-shot (RN/50) 17.3 19.2 16.0 21.9 18.6 2.6
CLIP B/32 (MoTIF) 56.3 44.9 51.3 61.0 53.4 6.9
Global CBM (B/32) 42.6 27.5 31.6 47.0 37.2 9.1
Zero-shot (B/32) 24.6 19.8 22.0 26.4 23.2 2.9
CLIP L/14 (MoTIF) 71.5 62.3 66.7 76.8 69.3 6.2
CLIP L/14 (MoTIF-ST) 73.2 64.8 65.0 80.8 73.5 12.2
Global CBM (L/14) 61.3 44.7 48.7 66.4 55.3 10.2
Zero-shot (L/14) 32.4 28.3 26.5 37.0 31.1 4.7
SigLIP L/14 (MoTIF) 76.1 62.1 73.1 82.7 73.5 8.6
Global CBM (SigLIP L/14) 59.9 44.7 53.3 70.6 57.1 10.9
Zero-shot (SigLIP L/14) 28.9 18.0 20.9 26.6 23.6 5.0
PE L/14 (MoTIF) 87.3 74.7 83.1 89.4 83.6 6.5
PE L/14 (MoTIF-ST) 86.2 74.7 82.3 91.2 83.6 7.0
Global CBM (PE L/14) 81.0 58.7 72.0 79.9 72.9 10.3
Zero-shot (PE L/14) 40.5 36.6 36.8 51.6 41.4 7.0
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 64.1 62.3 62.1 62.8 1.1
Global CBM (RN/50) 58.6 60.1 59.2 59.3 0.8
Zero-shot (RN/50) 29.3 30.1 30.1 29.8 0.5
CLIP B/32 (MoTIF) 65.9 66.8 63.3 65.3 1.8
Global CBM (B/32) 62.0 63.0 59.8 61.6 1.6
Zero-shot (B/32) 38.4 37.9 38.0 38.1 0.3
CLIP L/14 (MoTIF) 73.8 73.9 72.2 73.3 1.0
CLIP L/14 (MoTIF-ST) 75.5 75.3 73.6 74.8 1.0
Global CBM (L/14) 68.5 68.8 67.9 68.4 0.5
Zero-shot (L/14) 45.8 45.6 45.6 45.7 0.1
SigLIP L/14 (MoTIF) 74.8 74.4 70.4 73.2 2.4
Global CBM (SigLIP L/14) 66.3 66.0 62.6 65.0 2.1
Zero-shot (SigLIP L/14) 48.4 50.0 49.5 49.3 0.8
PE L/14 (MoTIF) 79.9 79.3 79.6 79.6 0.3
PE L/14 (MoTIF-ST) 79.9 80.0 78.8 79.6 0.7
Global CBM (PE L/14) 74.0 75.0 74.1 74.4 0.6
Zero-shot (PE L/14) 56.7 57.3 56.2 56.7 0.6
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 82.4 82.5 83.4 82.8 0.6
Global CBM (RN/50) 80.7 80.0 79.3 80.0 0.7
Zero-shot (RN/50) 56.5 56.9 58.3 57.2 0.9
CLIP B/32 (MoTIF) 84.4 85.7 86.7 85.6 1.2
Global CBM (B/32) 82.2 82.5 83.6 82.8 0.7
Zero-shot (B/32) 59.4 60.2 60.1 59.9 0.4
CLIP L/14 (MoTIF) 92.4 93.7 93.6 93.2 0.7
CLIP L/14 (MoTIF-ST) 92.9 94.4 94.0 93.8 0.8
Global CBM (L/14) 88.8 91.0 90.3 90.0 1.1
Zero-shot (L/14) 71.1 70.6 70.1 70.6 0.5
SigLIP L/14 (MoTIF) 93.3 93.8 94.9 94.0 0.8
Global CBM (SigLIP L/14) 90.0 91.0 90.5 90.5 0.5
Zero-shot (SigLIP L/14) 80.0 81.9 79.2 80.4 1.4
PE L/14 (MoTIF) 94.6 95.7 95.8 95.4 0.7
PE L/14 (MoTIF-ST) 95.7 96.4 96.9 96.3 0.6
Global CBM (PE L/14) 94.6 93.9 95.0 94.5 0.6
Zero-shot (PE L/14) 73.8 75.6 74.4 74.6 0.9
C.2 Performance on the Ablation Concept Set
We additionally evaluate all methods using the non-agentic concept sets generated via an LLM, as described in Section 2.1. The corresponding results are reported in Table 9. Overall, both Global CBM and MoTIF exhibit slightly reduced performance compared to the agentic setting, which we attribute to the smaller and less complete concept set, limiting the amount of task-relevant information captured in the bottleneck.
Table 9: Performance comparison (% Top-1 accuracy). Mean $\pm$ standard deviation on train-test splits on Breakfast Actions, HMDB51, UCF101, and SSv2 with different CLIP-based backbones with the non-agentic concept set. We report seconds per training epoch next to the accuracy scores for all MoTIF variants.
@@ C.2 Performance on the Ablation Concept Set
MoTIF (ours)
MoTIF (RN/50) 52.8 $\pm$ 6.9 (4.2) 62.8 $\pm$ 1.1 (0.9) 82.8 $\pm$ 0.6 (1.5) 16.0 (10.0)
MoTIF (ViT-B/32) 53.4 $\pm$ 6.9 (4.2) 65.3 $\pm$ 1.8 (0.9) 85.6 $\pm$ 1.2 (1.5) 17.5 (9.9)
MoTIF (ViT-L/14) 69.3 $\pm$ 6.2 (4.3) 73.3 $\pm$ 1.0 (0.8) 93.2 $\pm$ 0.7 (1.5) 20.4 (10.1)
MoTIF-ST (ViT-L/14) 71.1 $\pm$ 7.7 (7.2) 74.8 $\pm$ 1.0 (1.8) 93.8 $\pm$ 0.9 (3.3) 23.9 (26.8)
MoTIF (SigLIP-L/14) 73.5 $\pm$ 8.6 (4.2) 73.2 $\pm$ 2.4 (0.8) 94.0 $\pm$ 0.8 (1.5) 22.4 (9.8)
[…]
MoTIF-ST (PE-L/14) 84.1 $\pm$ 6.4 (7.3) 79.6 $\pm$ 0.7 (1.8) 96.3 $\pm$ 0.6 (3.3) 35.1 (26.7)
Non-interpretable video models
TSM (Lin et al., 2019) 59.111footnotemark: 1 73.5 95.9 61.7
No frame left behind (Liu et al., 2021) 62.011footnotemark: 1 73.411footnotemark: 1 96.411footnotemark: 1 62.711footnotemark: 1
@@
Appendix D Algorithms